OpenStudy (anonymous):

URGENT HELP: In ΔABC, if a = c = 16 centimeters and m∠B = 85°, what is b to two decimal places? I don't understand the Law of Sines when it comes to calculation...

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

is it like this ? 

OpenStudy (ikram002p):

if yes then m<A=m<C k ?

OpenStudy (anonymous):

I suppose, it doesn't come with a picture...

OpenStudy (ikram002p):

ohh then its like this :) then sin B/b =sin A /a=sin C/c ok ? just find m<A could u do that ? remember m<A=m<C

OpenStudy (anonymous):

Okay, give me one second...

OpenStudy (anonymous):

Okay, um...I don't know, how would to rewrite it in the formula? I'm sorry, this is the first time I ever use this. I'm homeschooled. :\

OpenStudy (anonymous):

Would it be something like sin85/16 for sinB/b or...? Because I was told to use the side opposite the angle.

OpenStudy (ikram002p):

well forget about sine law , angles sum of triangle is 180 right ? so let A snd C be x  so x+x+85=180 , solve for x

OpenStudy (anonymous):

47.5 for both x's?

OpenStudy (phi):

yes

OpenStudy (phi):

NOW you can do the Law of sines

OpenStudy (ikram002p):

ok nw u can use law of sine ^_^ since u have m<A=m<B =x

OpenStudy (anonymous):

Okay, but I don't understand out to do the Law of Sines...how would you plug in all the numbers?

OpenStudy (phi):

you need a "pair" a side and its opposite angle to form a ratio you then set the ratios equal to each other

OpenStudy (ikram002p):

ok so sin B/b =sin A /a then sin 80 /b = sin 47.5 /16 solve for b

OpenStudy (phi):

if you are looking for side b, then you need to use b! so therefore you need to use its opposite angle B that gives one ratio now pick another side/angle pair

OpenStudy (anonymous):

Okay, I am at a complete loss here...I wrote down here: sin47.5/b = sin85/16 = sin47.5/16 So far so good?

OpenStudy (phi):

you only use 2 out of the 3 ratios and you can be flexible about how to write the ratios (if you know what you are doing). But to stay "unconfused" for the time being write sin(Angle) divided by its opposite side